### v1

@@ A.2.1 Gradient Descent (GD)
A.2.2 Noisy Gradient Descent (NGD)
@@ A.2.6 SCRUB
${\small\widehat{\operatorname{\mathbb{E}}}_{(x,y)\sim S_{\textsf{train}}\setminus U}\left[\mathrm{KL}{\left(M_{\theta_{\textsf{initial}}}(x)\|M_{\theta}(x)\right)}+\ell(\theta;(x,y))\right]-\widehat{\operatorname{\mathbb{E}}}_{(x,y)\sim U}\left[\mathrm{KL}{\left(M_{\theta_{\textsf{initial}}}(x)\|M_{\theta}(x)\right)}\right]}$
@@ A.2.7 NegGrad+
Appendix B Additional Experiments
@@
  • Figures 10 and 9 show that changes in the size of the forget set do not qualitatively change conclusions.
@@
Figure 10: Varying the forgetset size for a GPT-2 (355M) trained on IMDb when using Gaussian poisons.
Appendix C Understanding Why Approximate Unlearning Fails?
C.1 Logistic Regression Experiment to Validate Hypothesis 1
@@
C.2 Linear Regression Experiment to Validate Hypothesis 2
@@
Next, we construct the poison set $S_{\textsf{poison}}$ for indiscriminate data poisoning attack discussed in Section 4.2, and by following the hyperparameters in Appendix A.1.3 (however, we only ran gradient canceling for 500 epochs). We generate 1000 poisoned samples that incur a parameter change with distance $\|\theta(S_{\textsf{corr}})-\theta(S_{\textsf{corr}}\setminus S_{\textsf{poison}})\|_{1}\approx 3.3$. We generate poisons with respect to 5 different initializations of the poison samples and report the averaged results in Figure 5(a).

### v2

@@ D.2.1 Gradient Descent (GD)
D.2.2 Noisy Gradient Descent (NGD)
@@ D.2.6 SCRUB
${\widehat{\operatorname{\mathbb{E}}}_{(x,y)\sim S_{\textsf{train}}\setminus U}\left[\mathrm{KL}{\left(M_{\theta_{\textsf{initial}}}(x)\|M_{\theta}(x)\right)}+\ell(\theta;(x,y))\right]-\widehat{\operatorname{\mathbb{E}}}_{(x,y)\sim U}\left[\mathrm{KL}{\left(M_{\theta_{\textsf{initial}}}(x)\|M_{\theta}(x)\right)}\right]}$
@@ D.2.7 NegGrad+
D.2.8 Selective Synaptic Dampening (SSD)
Selective Synaptic Dampening (SSD) was introduced in Foster et al. (2024) in order to unlearning certain forget set from a neural network without retraining it from scratch. SSD unlearns by dampening certain weights in the neural network which has a high influence on the fisher information metric corresponding to the forget set as compared to the remaining dataset. Given a model with weights $\theta$, suppose $I_{U}$ and $I_{S}$ denote the Fisher information matrix calculated over the forget set $U$ and the deletion set $S$ respectively. SSD performs unlearning by dampening the corresponding weights $\theta_{i}$ via the operation
$\theta_{i}=\begin{cases}\beta\theta_{i}&\text{if}\quad I_{U,i}>\alpha I_{S,i}\\ \theta_{i}&\text{if}\quad I_{U,i}\leq\alpha I_{S,i}\end{cases}$ (9)
where $i\in[\lvert\theta\rvert]$ and $I_{U,i}$ denotes the $i$th diagonal entry in the Fisher information matrix $I_{U}$. In the above, $\alpha$ is a selection-weighting hyperparameter, and
$\beta=\min_{i}\left\{\frac{\lambda I_{S,i}}{I_{U,i}},1\right\}$
for some hyperparameter $\lambda$.
Later works such as Schoepf et al. (2024a) explored a parameter-tuning-free variant of SSD. SSD has also been previously explored in terms of its ability to mitigate data poisoning. Goel et al. (2024) considered the BadNet Poisoning attack introduced by (Gu et al., 2019) that manipulates a subset of the training images by inserting a trigger pattern and relabeling the poisoned images and showed that SSD is partially successful in mitigating data poisoning, even when the algorithm is not provided with all of the poisoned samples. Building on this, Schoepf et al. (2024b) further evaluated SSD, and its variants, on two other data poisoning scenarios, (a) overlaying sin function on the base images, and (b) data poisoning by moving backdoor triggers, showing that SSD succeeds in mitigating data poisoning, hence arguing, that SSD is a reliable unlearning algorithm. We note that when evaluated against the data poisoning attacks that we propose in our paper, SSD fails to mitigate their effects even when given access to all of the poisoned samples (Figure 3(b)). The discrepancy between our observations and the prior works can be attributed to the nature of the data poisoning attacks considered, with our attacks being more adversarial in nature.
In our experiments, we implemented SSD using the open-source implementation provided by the authors Foster et al. (2024) for a diverse choice of hyperparameters, and none of them could mitigate the effects of data poisoning (see Figure 9).
[…]
We had some confusion between $\alpha$ vs $\frac{1}{\alpha}$. Is that fixed now?
Appendix E Experiments
E.1 Detailed comparison of different unlearning algorithms
While some methods outperform others, their effectiveness varies across different tasks. We mention our key observations below:
  1. $\bullet$
  2. $\bullet$
  3. $\bullet$
E.2 Comparison of Gaussian and Targeted Data Poisoning
Targeted data poisoning can be brittle. The targeted data poisoning attack (Geiping et al., 2021) is more brittle than our suggested Gaussian poisoning attack. Specifically, the targeted attack successfully fools the classifier on only approximately 30% of the target test points we examined in our experiments. Consequently, for approximately 70% of the test points, no unlearning analysis is possible. In our experimental setup, we randomly selected 100 target test points and applied the targeted poisoning attack. In 71.2% of these cases, the attack did not succeed, making it impossible to infer unlearning failure using these target test points.
E.3 Additional Experiments
@@
  • Figure 9 analyzes SSD performance across different hyperparameter settings.
  • Figure 10 shows that the standard MIA used in literature to evaluate unlearning efficacy is not a suitable measure for doing so.
@@
  • Figures 13 and 12 show that changes in the size of the forget set do not qualitatively change conclusions.
Figure 9: Analyzing the Performance of SSD Across Different Hyperparameters. We use a ResNet-18 model trained on Gaussian-poisoned CIFAR-10 training data. As the strength of the selection weighting parameter increases, SSD progressively scales more weights until the updated model’s performance begins to degrade. Notably, even under Gaussian data poisoning evaluation, a significant portion of the forget set remains identified as belonging to the trained model, indicating that SSD does not effectively facilitate unlearning.
(a) Standard unlearning MIA (Shokri et al., 2017)
(b) Using our suggested Gaussian poisons
Figure 10: The standard unlearning MIA is not a suitable test of unlearning efficacy – full tradeoff curves comparison. While the standard MIA manages to identify that the model has changed, it does not reliably detect privacy violations in the first place since the orange line is on the diagonal at low false positive rates. Here we have unlearned 1.5% of the data from a Resnet-18 trained on CIFAR-10. The unlearning was done using NGD with a noise level of $1e^{-7}$.
@@
Figure 13: Varying the forget size for a GPT-2 (355M) trained on IMDb with Gaussian poisons.
Appendix F Understanding Why Approximate Unlearning Fails?
F.1 Logistic Regression Experiment to Validate Hypothesis 1
@@
F.2 Linear Regression Experiment to Validate Hypothesis 2
@@
Next, we construct the poison set $S_{\textsf{poison}}$ for indiscriminate data poisoning attack discussed in Section 4.2, and by following the hyperparameters in Appendix D.1.3 (however, we only ran gradient canceling for 500 epochs). We generate 1000 poisoned samples that incur a parameter change with distance $\|\theta(S_{\textsf{corr}})-\theta(S_{\textsf{corr}}\setminus S_{\textsf{poison}})\|_{1}\approx 3.3$. We generate poisons with respect to 5 different initializations of the poison samples and report the averaged results in Figure 14(a).